jhonyy9 (jhonyy9):

how many same numbers may there exist by 4 digits multiplied by a 5th different number will result the first number inversed order abcd*e=dcba a,b,c,d and e are natural numbers

jhonyy9 (jhonyy9):

i have got one 2178*4 -------- 8712 there may exist ?

jhonyy9 (jhonyy9):

@Hero @Preetha @agent0smith

jhonyy9 (jhonyy9):

@ganeshie8 @mathmath333 @mathmale @Directrix

jhonyy9 (jhonyy9):

@TheSmartOne

OpenStudy (kainui):

Does this count: 0009*1000=9000

jhonyy9 (jhonyy9):

every digits need being different from zero till 9 - this is why i noted there abcd*e - different letters for different digits

jhonyy9 (jhonyy9):

@TheSmartOne ty

jhonyy9 (jhonyy9):

nice to meet you dear my friend

jhonyy9 (jhonyy9):

@Directrix

OpenStudy (mathmath333):

*

OpenStudy (mathmate):

Let the number (in general) be pqrs =\(1000p+100q+10r+s\) the number be multiplied by an integer k, then we want \(k(1000p+100q+10r+s)=1000s+100r+10q+p\) Solve for k \(\large k=\frac{1000s+100r+10q+p}{1000p+100q+10r+s}\) Set p=2, q=1, r=7, s=8 and evaluate k. Hint: k is the square of the first prime number.

OpenStudy (mathmate):

Oh, you want other ones, give me a sec!

OpenStudy (mathmate):

@jhonyy9 Here's what I've got for all 4 digit numbers (i.e. first digit is 1-9): k=1 for all palindromes, example 1*1001=1001, 1*2442=2442, etc. If k≠1, which is probably what you want, you have hit half of the lot: {2178, 1089} are the non-palindrome solutions ( k≠1 ) Someone brilliant can perhaps give the mathematical solution. If all digits including k must be different, your solution is the only one.

jhonyy9 (jhonyy9):

ty @mathmale yes sure how i wrote above all digits must be different so than this is sure that my solution is the only one ? ty.

OpenStudy (mathmate):

Exactly. You've obtained the one and only one solution where all five digits are different...by exhaustive search. :)

jhonyy9 (jhonyy9):

@BOOKISHFAIRY73

jhonyy9 (jhonyy9):

@bookishfairy73